OpenStudy (anonymous):

√48x^6=I need help doing this please.

OpenStudy (anonymous):

what are the factors of 48?

OpenStudy (anonymous):

1,2,3,4,6,8,12,16,24,48

OpenStudy (anonymous):

yes, I'm sorry, I meant can you factor it? 2,2,2,2,3 isn't it?

OpenStudy (anonymous):

so you have the square root of (2 2 2 2 3 x x x x x x) right?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

\[\sqrt{48x^6}=\sqrt{2^4*3}\sqrt{x^6}=4x^3\sqrt{3}\]

OpenStudy (anonymous):

so pull out as many pairs as you can, and see what is left

OpenStudy (anonymous):

animal is right, but I think he used magic

OpenStudy (anonymous):

so i got two sets of 2s and i set of 3xs. so that is 4(x^3)?

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

okay. thank you for your help.

OpenStudy (anonymous):

but dont forget what is left over stays under the radical

OpenStudy (anonymous):

so the 3 goes under the radical?

OpenStudy (anonymous):

can you help me with another one? a cube root?

OpenStudy (anonymous):

yep, can't pull it out so it hangs around out there